OpenStudy (anonymous):

Find the condition that the constants a and b must satisfy in order that y = ax + b satisfies the equation: x(dy/dx)^2 - y(dy/dx) = 2 Find the two solutions of this equation satisfying y = 3 when x = 2. Would appreciate a walk through as I'm working through this for revision, esp. at the initial re-arranging stage.

OpenStudy (amistre64):

a y'^2 - y y' = 2 is the question? this is differential equation stuff then right?

OpenStudy (anonymous):

Yes, x y'^2 - y y' = 2

OpenStudy (amistre64):

cause it could also be quadratic stuff

OpenStudy (anonymous):

differential stuff

OpenStudy (amistre64):

x(y'^2) -y(y') -2 = 0

OpenStudy (amistre64):

if y=3 and x=2 we get what: 2.-2 = -4 4 - 1 = 3 (y' +4)(y'-1) = 0 is what i come up with

OpenStudy (amistre64):

opps forgot to divide out the 2 :)

OpenStudy (amistre64):

(y' +2) (2y'-1) = 0

OpenStudy (amistre64):

... but i got know idea if this is helpful

OpenStudy (anonymous):

Those are the solutions, but how can I find the constraints?

OpenStudy (anonymous):

\[\frac{dy}{dx} = \frac{d}{dx}(ax + b) = \frac{d}{dx}ax+\frac{d}{dx}b=a+0=a.\] \begin{eqnarray*}x \left(\frac{dy}{dx}\right)^2 - y\left(\frac{dy}{dx}\right) &=&xa^2 - (ax + b)a \\&=&xa^2 - a^2x-ab\\&=&-ab=2.\end{eqnarray*} If \[y(3) = 2,\]\[y(3) = 3a + b = 2 \Rightarrow b = 2-3a \Rightarrow a(2-3a) = -2\] \[\Rightarrow 3a^2 -2a - 2 = 0 \Rightarrow a = \frac{1}{3}\left(1 \pm \sqrt{7}\right) \land b = 1 \mp \sqrt{7}.\]

OpenStudy (amistre64):

gonna have to rely on one of them there smarter than mes for that one ;)

OpenStudy (anonymous):

so in general should I sub in a for y'

OpenStudy (anonymous):

No, that's only because the problem said that\[y = ax+b.\]If we didn't have that information, the problem would be much more difficult.

OpenStudy (anonymous):

Thanks for the help, is there any way of saving threads? Just found this openstudy thing this evening, really going to help with my revision.

OpenStudy (amistre64):

your questions are saved on your profile; just click on your screenname

OpenStudy (anonymous):

Sweet, thanks